[Brahms-l] Shift report 20050622 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Wed Jun 22 2005 - 08:01:48 EDT
Shifter leader was SB
Shifters were 


<br>
1. Beam from 1:30 AM continues into next shift<br>
   Run 15739 going on<br>
2. Runs 15729,15731,15734-15738<br>
Trig2(478k),Trig3(175k),Trig4(118k),Trig5(689k)<br>

3. No detector problems. IoServer crashed once and the data for<br>
 run 15730 seems lost. Runs 15726,15727,15728 should not be used.<br>
Detectors were accidentally off.<br>
-----------------------------------------------<br>
Time           Blue Pol. (%)     Yellow Pol.(%)<br>
-----------------------------------------------<br>
At injection  -46.11 +/- 2.43  -51.45 +/- 2.29<br>
4:28 AM       -42.62 +/- 2.40  -58.20 +/- 2.27<br>
7:28 AM       -39.78 +/- 2.37  -53.16 +/- 2.28<br>


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed Jun 22 08:02:06 2005

This archive was generated by hypermail 2.1.8 : Wed Jun 22 2005 - 08:02:18 EDT